nolegirl24:

Which of the following is the graph of f(x) = x2 − 4x − 5? graph of a quadratic function with a minimum at 2, negative 9 and x intercepts at negative 1 and 5 graph of a quadratic function with a minimum at 3, negative 4 and x intercepts at 1 and 5 graph of a quadratic function with a minimum at 2.5, negative 2.4 and x intercepts at 1 and 4 graph of a quadratic function with a minimum at negative 1.5, negative 6.2 and x intercepts at 1 and negative 4 (IF YOU NEED A PIC I WILL TRY!!!!)

dude:

Yeah uh I might need a picture

nolegirl24:

Okay one sec

dude:

To find where it intercepts the x axis you can factor this \(f(x) = x^2 − 4x − 5\) What are the multiples of 5?

nolegirl24:

1,5 5,1

dude:

Since 5 is negative you need a positive and a negative number \((x+number) (x-number)\) Which order of these numbers can add up to 4?

dude:

-4*

nolegirl24:

-5,1?

dude:

Right So we have \((x-5)(x+1)\) To find the x-intercept (When the graph touches the x-axis) we set them equal to 0 \(x-5=0\) and \(x+1=0\)

nolegirl24:

Alright

nolegirl24:

So it would be 5,-1?

dude:

Yes, so x should be 5 and -1 when it touches the x axis

nolegirl24:

So graph A?

dude:

Yes

nolegirl24:

thanksssssssssss

dude:

Sure no problem!

nolegirl24:

You are one of the biggest help ever!